OpenStudy (anonymous):

Which sentence correctly uses a nominative case pronoun? A. The girl on the left is her. B. Did Isaac and him find the missing game controller? C. Ralph and them are waiting for us at the Metro station. D. It was Max and I on the trampoline.

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

D

OpenStudy (anonymous):

Thanks that is it are you good at math???

OpenStudy (blurbendy):

pretty good. you're catching up with me though, so I'll see if I can still help. you can just post here if you want.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

well about to eat lunch how about after my lunch about 2:00 o clock

OpenStudy (blurbendy):

sure, I'll stay on

OpenStudy (anonymous):

ok r u ready:)

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

A coffee pot makes 6 cups of coffee in 1 1/2min. How many cups of coffee can the coffee pot make in 1 min? cups

OpenStudy (blurbendy):

90 seconds = 6 cups of coffee 90/6 = 15 1 cup = 15 seconds 15 + 15 + 15+ 15 = 60 seconds = 1 min = 4 cups of coffee

OpenStudy (anonymous):

A coffee pot makes 6 cups of coffee in 1 1/2min. How many cups of coffee can the coffee pot make in 1 min? cups

OpenStudy (blurbendy):

i answered

OpenStudy (anonymous):

ok i have another one please

OpenStudy (anonymous):

Dora can run 3/5 mi in 3/20 h. How far can she run in 1 h? mi

OpenStudy (blurbendy):

are there answer choices?

OpenStudy (anonymous):

no

OpenStudy (blurbendy):

Sorry, I got distracted Anyways 3/20 = 9 minutes 60/9 = 6 and 2/3 6 and 2/3 * 3/5 = 4 4 miles

OpenStudy (blurbendy):

@cjackson2002